OpenStudy (dls):

A ball is dropped from a height of 90m on a floor.At each collision with the floor ,the ball loses one tenth of its speed.Plot the speed time graph of its motion between t=0 to 12s

OpenStudy (anonymous):

S= ut + 1/2 * a* t^2

OpenStudy (dls):

t=4.29 s

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

nw find final velocity

OpenStudy (anonymous):

v = u + at

OpenStudy (dls):

42.04m/s

OpenStudy (dls):

9/10v=37.84m/s

OpenStudy (anonymous):

yes..... procede

OpenStudy (dls):

i could just not the plot the graph

OpenStudy (anonymous):

The graph would look something like this:

OpenStudy (anonymous):

The graph will be discontinuous since at the time the ball hits the ground its speed drops abruptly and not gradually.

OpenStudy (anonymous):

The picture shows the discontinuities joined.If you watch closer the graph in the picture shows at time t=4 seconds the particle may have any speed between 42 and 32.But that is meaningless.So it is actually discontinuous.

OpenStudy (anonymous):

 yea i agree with @kartiksriramk ...it should be this way.